Solve the system by substitution.
5x - 2y = 10
y = 3x - 5

Found 2 solutions by checkley75, mommyengelman1:
Answer by checkley75(3666)

Y=3X-5 NOW REPLACE Y BY (3X-5) IN THE OTHER EQUATION & SOLVE FOR X.
5X-2(5X-5)=10
5X-10X+10=10
-5X=10-10
5X=0
X=0 ANSWER.
Y=3*0-5
Y=-5 ANSWER.
PROOF
5*0+2*5=10
0+10=10
10=10

Answer by mommyengelman1(2)

original
5x-2y=10
y=3x-5
5x-2(3x-5)=10
Remember order of operations and combine like terms.
5x-6x+10=10
-10 -10
5x-6x=0
-x=0
/-1 /-1
x=0
To find y we then subtitute our found value for x it into the original problem to complete our pair.
5(0)-2y=10
0-2y=10
-2y=10
/-2 /-2
y=-5
our ordered pair is (0,-5)
To check to see if our answers are correct, then we plug in our answers for our ordered pair into the original equation.
5(0)-2(-5)=10
0+10=10
10=10
true